# Samuele Sabella Istituto di Informatica e Telematica Consiglio Nazionale delle Ricerche Pisa, Italy

### v1

@@ I Introduction
The Federated Learning paradigm leverages the use of multiple entities connected within a network, empowered with enough computational power to solve the task of knowledge extraction and model training. Each device extracts knowledge from locally available data by a training machine learning model and shares it with its neighbors, resulting in local knowledge spread across the network without the raw data leaving the device. Thus, problems that typically require access to all the data produced by the system, as done in centralized solutions, are solved using multiple learning devices trained in parallel on locally available data. Federated learning is studied within environments where it is not possible or convenient to share raw data. Healthcare facilities have strict privacy-related problems preventing them from sharing patients’ data outside the infrastructure where they are collected or produced. Sharing locally trained neural networks has been one key solution to combine the knowledge coming from multiple sources and improve the predictions’ performance [1][2]. Internet of Things (IoT) networks usually have bandwidth limits that prevent sharing large volumes of data across the network. Sharing a compressed data representation may represent the only solution to learning a shared objective between devices using all the data available to the system. When computing power is not scarce and learning is required, decentralized learning could be a valid approach to avoid sending large quantities of data to a central server where learning is done via a centralized technique. Moreover, end users are increasingly concerned about privacy. Sharing only the knowledge extracted from the end devices without exposing the data used to extract that knowledge offers an effective alternative to centralized learning solutions that require all data to be accessible during training.
@@ II-B Attacks against decentralized learning
figureThe hub’s local dataset undergoes corruption at $75\%$ with the corruption strength $\alpha=0.95$. The corrupted images are built using the en
@@ III-A Decentralized learning
At the end of the local training phase, device $i$ transmits $\tilde{\mathbf{w}}_{i}$ to its neighboring nodes. The model for the next communication round is then updated by aggregating the received models in a layer-wise manner, weighted by the respective dataset sizes. Given $\mathcal{N}(i)$, the set containing node $i$ and its neighbors, the model update at communication round $t$ is defined as:
@@ IV-B Dataset and data distribution
Figure 1: Examples of interpolated images of nine with images of class four with $\alpha=0.5$ (left) and $\alpha=0.95$ (right)
@@ V Results
Before proceeding with a detailed analysis, we we first highlight a key observation: our experiments indicate that decentralized federated averaging is largely robust to the presence of corrupted data in the system. Specifically, we found that data corruption effects become clearly pronounced when its strength is set to $\alpha=0.95$. Consequently, unless stated otherwise, all results presented in this work have been obtained using this corruption strength. Refer to Figure 1 for a qualitative illustration of the impact of corruption for different $\alpha$.
@@
TABLE II: Comparison of final F1 scores w/o corruption and w/ corruption ($\alpha=0.95$ and $p=0.9$)
@@ VI Discussion and Conclusion
n this work, we investigated the impact of corrupted data on decentralized federated learning (DFL) and compared its resilience to federated learning (FL) and centralized learning. Our evaluation was conducted on a realistic communication network modeled as a heavy-tailed Barabási–Albert (BA) graph, with a focus on a standard image classification task. We simulated decentralized environments where some nodes’ local datasets contained low-quality samples, mimicking real-world scenarios in which data degradation arises from sensor noise, faulty data augmentation, or intentional adversarial interference. To introduce controlled corruption, we leveraged interpolations in the latent space of a pre-trained GAN, enabling targeted corruption of specific labels with varying intensities. Our experiments examined both balanced and unbalanced corruption distributions, considering scenarios where corrupted data was either spread across multiple nodes or concentrated on a single node, including highly connected hubs. We also explored different corruption levels, ranging from minor perturbations to extreme cases where up to 90% of class 9 samples were corrupted.

### v2

@@ I Introduction
The Federated Learning paradigm leverages the use of multiple entities connected within a network, empowered with enough computational power to solve the task of knowledge extraction and model training. Each device extracts knowledge from locally available data by training a machine learning model and shares it with its neighbors, resulting in local knowledge spread across the network without the raw data leaving the device. Thus, problems that typically require access to all the data produced by the system, as done in centralized solutions, are solved using multiple learning devices trained in parallel on locally available data. Federated learning is studied within environments where it is not possible or convenient to share raw data. Healthcare facilities have strict privacy-related problems preventing them from sharing patients’ data outside the infrastructure where they are collected or produced. Sharing locally trained neural networks has been one key solution to combine the knowledge coming from multiple sources and improve the predictions’ performance [1][2]. Internet of Things (IoT) networks usually have bandwidth limits that prevent sharing large volumes of data across the network. Sharing a compressed data representation may represent the only solution to learn a shared objective between devices using all the data available to the system. When computing power is not scarce and learning is required, decentralized learning could be a valid approach to avoid sending large quantities of data to a central server where learning is done via a centralized technique. Moreover, end users are increasingly concerned about privacy. Sharing only the knowledge extracted from the end devices without exposing the data used to extract that knowledge offers an effective alternative to centralized learning solutions that require all data to be accessible during training.
@@ II-B Attacks against decentralized learning
figureThe hub’s local dataset undergoes $75\%$ corruption with a corruption strength of $\alpha=0.95$. The corrupted images are generated by interpolating samples from class four with samples from class nine using a pretrained GAN, where the interpolation is controlled by the parameter $\alpha$.
@@ III-A Decentralized learning
At the end of the local training phase, the device $i$ transmits $\tilde{\mathbf{w}}_{i}$ to its neighboring nodes. The model for the next communication round is then updated by aggregating the received models in a layer-wise manner, weighted by the respective dataset sizes. Given $\mathcal{N}(i)$, the set containing node $i$ and its neighbors, the model update at communication round $t$ is defined as:
@@ IV-B Dataset and data distribution
Figure 1: Examples of interpolated images of nine with images of class four with $\alpha=0.5$ (left) and $\alpha=0.95$ (right).
@@ V Results
Before proceeding with a detailed analysis, we first highlight a key observation: our experiments indicate that decentralized federated averaging is largely robust to the presence of corrupted data in the system. Specifically, we found that data corruption effects become clearly pronounced when its strength is set to $\alpha=0.95$. Consequently, unless stated otherwise, all results presented in this work have been obtained using this corruption strength. Refer to Figure 1 for a qualitative illustration of the impact of corruption for different $\alpha$.
@@
TABLE II: Comparison of final F1 scores w/o corruption and w/ corruption ($\alpha=0.95$ and $p=0.9$).
@@ VI Discussion and Conclusion
In this work, we investigated the impact of corrupted data on decentralized federated learning (DFL) and compared its resilience to federated learning (FL) and centralized learning. Our evaluation was conducted on a realistic communication network modeled as a heavy-tailed Barabási–Albert (BA) graph, with a focus on a standard image classification task. We simulated decentralized environments where some nodes’ local datasets contained low-quality samples, mimicking real-world scenarios in which data degradation arises from sensor noise, faulty data augmentation, or intentional adversarial interference. To introduce controlled corruption, we leveraged interpolations in the latent space of a pre-trained GAN, enabling targeted corruption of specific labels with varying intensities. Our experiments examined both balanced and unbalanced corruption distributions, considering scenarios where corrupted data was either spread across multiple nodes or concentrated on a single node, including highly connected hubs. We also explored different corruption levels, ranging from minor perturbations to extreme cases where up to 90% of class 9 samples were corrupted.